OpenStudy (anonymous):

Rework problem 1 in section 4.1 of your text, involving the flipping of a loaded coin, but assume that Pr[H] = 0.3. Also, assume that the coin is flipped 4 times, and the random variable X is defined to be 3 times the number of heads minus 2 times the number of tails. How many different values are possible for the random variable X?

OpenStudy (anonymous):

grind it till you find it no heads : -8 one head: 3-6=-3 two heads: 6-4=2 three heads: 9-2=7 four heads: 12